even though we have our books in algebra...
I do not understand what our teacher is tackling about... maybe he is so fast in discussing of the topics so that we can't relate to him...
In our homework... i don't know what i am going to do because this topic is not yet discuss unto us but i want to learned it so i can get high score during our quizzes...i just want to know if how come the answer goes like this...an=2n-6?
our problem or exercise needed to solve is that:
write a formula for the nth term of arithmetic sequence...
==> -4,-2,0,2,4,...
can you help me please... can you explain it to me?
it is not from the book... it's from our math teacher..

Answer by stanbon(75887)

write a formula for the nth term of arithmetic sequence...
==> -4,-2,0,2,4,...
------------------------
The formula for the nth term of an arithmetic sequence is
a(n) = a(1)+(n-1)d
a(1) is the 1st term of the sequence
d is the difference between any term and the one before it.
-------------------
Your Problem
The 1st term is -4
d = 2-0 = 2
So the nth term is a(n)=-4+(n-1)*2
a(n) = -4+2n-2
a(n) = -6+2n
==================
Cheers,
Stan H.